OpenStudy (horsegal244):

What is the volume of a 75 cubic inches rectangular prism if scaled down by 1/10? I will give medals and fan!

OpenStudy (horsegal244):

@brunomcv

OpenStudy (horsegal244):

@jim_thompson5910

OpenStudy (horsegal244):

@SolomonZelman

jimthompson5910 (jim_thompson5910):

Each dimension is scaled by 1/10?

OpenStudy (horsegal244):

Yea

OpenStudy (horsegal244):

@jim_thompson5910

OpenStudy (amistre64):

how do we find the volume to start with?

OpenStudy (horsegal244):

Height x width x length

OpenStudy (amistre64):

good, so 75 = lwh now what happens when we divide each linear measure by 10? V = l/10 * w/10 * h/10

OpenStudy (horsegal244):

I don't Know Im sorry

OpenStudy (amistre64):

\[V=\frac{l}{10}\frac{w}{10}\frac{h}{10}=\frac{lwh}{10^3}\] but we are told what lwh equals already, its 75 so what is the value of the new Volume?

OpenStudy (anonymous):

each of three dimensions gets scaled down by \(1/10\), so the volume scales by \((1/10)^3=1/10^3=\dots\)

OpenStudy (amistre64):

if its scaled down by 1/10, doesnt that mean that we have 9/10 to deal with in the volume?

OpenStudy (horsegal244):

L 7 1/2 over 10 w 4 over 10 h 2 1/2 over 10

OpenStudy (amistre64):

the actual numbers for l,h, and w are irrelevant. lwh = 75 if we scale down by 1/10 ... thats the trouble spot to read thru. it reads to me as: (1 - 1/10) is our scale factor. so instead of altering it by (.10)^3 we would be altering it by (.90)^3 the question is, how correct is my thinking?

OpenStudy (horsegal244):

I don't really know can u please just solve it for me....

OpenStudy (amistre64):

no, i cant just solve it for you.

OpenStudy (horsegal244):

Please I'm begging you i am horrid at math its not funny

OpenStudy (amistre64):

asking or begging me for a direct answer would simply force me to suspend your account.

OpenStudy (horsegal244):

Well thanks